OpenStudy (anonymous):

Which of the following represents the vertex and x-intercepts of the function given below? y = x^2 + 8x + 12 A. Vertex: (2,6) Intercepts: x = 6, 2 B. Vertex: (4,60) Intercepts: x = 6, -2 C. Vertex: (-4,-4) Intercepts: x = -6, -2 D. Vertex: (4,-4) Intercepts: x = -4, -3

OpenStudy (anonymous):

correct me if im wrong but I got C!

OpenStudy (kenljw):

y = x^2 + 8x + 12 y = (x + 4)^2 -16 +12 y = (x + 4)^2 - 4 Vertex (-4, -4) x -intercept is when y = 0, (-2, 0), (6, 0) y - intercept is when x = 0, (0, 12)